OpenStudy (joannablackwelder):

Use Coulomb's law to calculate the ionization energy in kJ/mol of an atom composed of a proton and an electron separated by 151.00 pm .What wavelength of light would have sufficient energy to ionize the atom?

OpenStudy (joannablackwelder):

I used Coulomb's law to get the force of attraction, which I got to be about 1.01E-8N, but I'm not sure how to find the Joules from there. Since to ionize, we are essentially moving the electron and the proton and "infinite" distance apart.

OpenStudy (joannablackwelder):

@abb0t

OpenStudy (joannablackwelder):

@Photon336

OpenStudy (photon336):

@joannablackwelder interesting question \[\frac{ q_{1}q_{2}k }{ r^{2} } = F_{c}\] q_{1} = charge proton q_{2} = charge of electron r = {distance between them} \[E = \frac{ hc }{ \lambda }\] \[\lambda = \frac{ hc }{ E }\] @kainui

OpenStudy (joannablackwelder):

Thanks, @Photon336 , but how do you get from F to E?

OpenStudy (joannablackwelder):

without a distance of separation that they would need to be to be sufficiently away from each other?

OpenStudy (photon336):

Yeah, i'm thinking sufficient distance but i'm not sure how to find it. You know who might be able to help @Michele_Laino but she's not on

OpenStudy (joannablackwelder):

Thanks :-)

OpenStudy (joannablackwelder):

@chmvijay @aaronq @Cuanchi

OpenStudy (joannablackwelder):

@Kainui

OpenStudy (joannablackwelder):

@TheSmartOne

OpenStudy (aaronq):

Pretty sure you just find the integral of the force versus distance (radius), which is "work", and integrate from 151 pm to infinity.

OpenStudy (joannablackwelder):

Ah!!! Thanks!

OpenStudy (aaronq):

no problem!